OpenStudy (anonymous):

27. Find the 20th term of the arithmetic sequence 9, 13, 17, 21, ... 28. Find the 17th term of the arithmetic sequence 12, 3, –6, –15, ...

OpenStudy (haseeb96):

A=9, d is the difference d=13-9=5 N=20 The formula of arthematic sequence is a+(N-1).d Now just put all those values and show me your answer

OpenStudy (anonymous):

855? @Haseeb96

OpenStudy (anonymous):

a=9 d=13-9=4 n=20 \[t _{n}=a+(n-1)d\]